# How Many Van Goghs Does It Take to Van Gogh? Finding the Imitation Threshold

### v1

@@ D.1 Collection of Reference Images for Human Faces Domain
    Candidate set: First, we retrieved the first hundred images by searching a person’s name on Google Images. We used SerpAPI SerpApi (2024) as a wrapper perform the searches.
@@
    One strategy is for the faces to form a graph where the vertices are the face embeddings and the edges connecting two embeddings have a weight equal to the cosine similarity between them, and we select a dense k-subgraph (Lanciano et al., 2024) from this graph. Selecting such a subgraph means finding a mutually homogeneous subset. We can find the vertices of this dense k-subgraph by cardinality-constrained submodular function minimization (Bilmes, 2022; Nagano et al., 2011) on a facility location function (Bilmes, 2022). We run this minimization and select a subset of images (at least of size ten) that has the highest average cosine similarity between each pair of images.
@@
  $\triangleright$ Retrieve initial image set using SerpAPI
3
@@
  $\triangleright$ Select candidate set using submodular minimization
4
@@
Appendix E Implementation Details of MIMETIC2 for Human Face Imitation
E.1 Filtering of Training Images
Images whose captions mention the concept of interest often do not contain it (as shown with Mary-Lee Pfeiffer in Figure 4). As such, we filter images where the concept does not appear in the image, which we detect using a dedicated classifier. In what follows we describe the filtering mechanism.
@@ E.1 Filtering of Training Images
We collect reference images for each person using SerpAPI as described in Appendix D. These images are the gold standard images that we manually vet to ensure that they contain the target person of interest (see Appendix D for the details). We use the reference images to filter out the images in the pretraining dataset that are not of this person. Concretely, for each person we use a face embedding model Deng et al. (2020) to measure the similarity between the faces in the reference images and the faces in the images from the pretraining datasets whose captions mention this person. If the similarity of a face in the pretraining images to any of the faces in the reference images is above a certain threshold, that face is considered to belong to the person of interest. We determine this threshold to distinguish faces of the same person from faces of different persons in the next paragraph. Note that this procedure already filter outs any image that does not contain a face, because the face embedding model would only embed an image if it detects a face in that image.
@@
Determining Filtering Threshold: The next step is to determine the threshold for which we consider two faces to belong to the same person. For this purpose, we measure the similarity between pairs of faces of the same person and the similarity between pairs of faces of different persons Since the reference images for each person is manually vetted to be correct, we use these images for this procedure. We plot the histogram of the average similarity between the faces of the same person (blue colored) and the similarity between faces of different persons (red colored) in Figure 7. We see that the two histograms are well separated, with the lowest similarity value between the faces of the same person being 0.56 and the highest similarity value between the faces of different persons being 0.36. Therefore any threshold value between 0.36 and 0.56 can separate two face of the same person, from the faces of different people. In our experiments, we use the midpoint threshold of 0.46 (true positive rate (tpr) of 100%; false positive rate (fpr) of 0%) to filter any face in the pretraining images that do not belong to the person of interest. The filtering process gives us both the image frequency a person in the pretraining data, and the pretraining images that we compare the faces in the generated images to measure the imitation score.
E.2 Measurement of Imitation Score
@@
To measure the imitation between the training and generated images of a person, we compute the cosine similarity between the face embeddings of the faces in their generated images and their filtered training images from the previous step. However, measuring the similarity using all the pretraining images can underestimate the actual imitation. This is because several individuals have significant variations in their faces in the pretraining images and the text-to-image model does not capture all these variations. For example, consider the pretraining images of Samuel L. Jackson in Figure 8(a). These images have significant variations in beard, hair, and age. However, when the text-to-image model is prompted to generate images of Samuel L. Jackson, the generated images in Figure 8(b) only show a specific facial characteristic of him (middle-aged, bald, with no or little beard). Since MIMETIC2’s goal is not to measure if a text-to-image model captures all the variations of a person, we want to reward the model even if it has only captured a particular characteristic (which it has in this case of Samuel L. Jackson). Therefore, instead of comparing the similarity of generated images to all the training images, we compare the similarity to only the ten training images that have the highest cosine similarity to the generated images on average.
Appendix F Implementation Details of MIMETIC2 for Art Style Imitation
F.1 Filtering of Training Images
@@
Filtering Non-Art Images: To filter non-art images from the pretraining dataset, we use a classifier that separates art images from non-art images. Concretely, we embed the pretraining images using a CLIP ViT-H/14 Ilharco et al. (2021) image encoder and measure the cosine similarity of the image embeddings and the text embeddings of the string ‘an artwork’, embedded using the text encoder of the same model. Only when the similarity between the embeddings is higher than a threshold described below, we consider those pretraining images as an artwork. To determine this threshold, we choose a similarity score that separates art images from non-art images. We use the images from the Wikiarts dataset Saleh & Elgammal (2016) as the (positive) art images and MS COCO dataset images Lin et al. (2014) as the (negative) non-art images. Note that MS COCO dataset was collected by photographing everyday objects that art was not part of, making it a valid set of negative examples of art.
We plot the histogram of cosine similarity of the embeddings of art and non-art images to the text embedding of ‘an artwork’ (see Figures 9(a) and 9(b). We observe that the art and non-art images both the artist groups are well separated (although not perfect, Figure 10 and Figure 11 shows examples of misclassified and correctly classfied images from both datasets). We choose the threshold that maximizes the F1 score of the separation (0.182 for the classical artists and 0.177 for the modern artists).
@@
Figure 10: Images that are misclassified by our art vs. non-art threshold in Figure 9(a).
@@
Figure 11: Images that are correctly classified by our art vs. non-art threshold in Figure 9(a).
@@
Filtering Images of Other Art Styles: Similar to the case of human faces, not all art images whose captions mention an artist were created by that artist. We want to filter out such images. For this purpose, we collect reference images for each artist (see Appendix D for details) and use them to classify the training images that belong to the artist of interest. Concretely, we measure the similarity between the pretraining images and the reference images of each artist, and only retain images whose similarity to the reference images is higher than a threshold.
To determine this threshold, we measure the similarity between pairs of art images of the same artists and pairs of art images from different artists. We embed the images using an art style embedding model Somepalli et al. (2024) and plot the histogram of similarities between art images of the same artist (blue colored) and art images of different artists (red colored) in Figure 12(a) for classical artists and Figure 12(b) for modern artists. We see that the two histograms are well separated (although not perfect, Figure 13 shows paintings by two artists whose art style is very similar and cannot be distinguished by our threshold). We choose the threshold that maximizes the F1 score of the separation between these two groups (0.278 for classical artists and 0.288 for modern artists). The retained images give us both the image counts of each artist and the training images that we compare to the generated images to measure the imitation score.
@@ F.2 Similarity Measurement
We embed all the generated images and the filtered pretraining images using the art style embedding model (Somepalli et al., 2024) and measure the cosine similarity between each pair of generated and pretraining images. Similar to the case of the human faces, we do not want to underestimate the art style similarity between the generated and training images by comparing the generated images to all the training images of this artist. Therefore, we measure the similarity of generated images to the ten training images that are on average the most similar to the generated images.
Appendix G Imitation Thresholds of SD models in Series 1 and 2
@@
Our experimental results in Section 6 found that for most domains the imitation thresholds for SD1.1 and SD1.5 are almost the same, while being higher for SD2.1. We hypothesized that the difference is due to their different text encoders. All models in SD1 series use the same text encoder from CLIP, whereas SD2.1 uses the text encoder from OpenCLIP. To test the validity of this hypothesis, we repeated the experiments for all models in SD1 series for politicians and computed their imitation thresholds. Table 6 shows the thresholds for the politicians. We find that the imitation thresholds for all the models in SD1 series is almost the same, and is lower than the threshold for SD2.1 model. This evidence supports our hypothesis of the difference in the text-encoders being the main reason for the difference in the imitation thresholds.
@@
Appendix H Change Points
Table 7 we show all the change points that PELT found for each experiment (Table 3 reports the first change point as the imitation threshold).
Table 7: Imitation Thresholds for human face and art style imitation for the different text-to-image models and datasets we experiment with.
@@ Appendix H Change Points
Appendix I All Results: The Imitation Threshold
@@
In this section, we estimate the imitation threshold for human face and art style imitation for three different text-to-image models. Figure 17, Figure 18, and Figure 19 show the image counts of celebrities on the x-axis (sorted in increasing order of image counts) and the imitation score of their generated images (averaged over the five image generation prompts) on the y-axis. The images were generated using SD1.1, SD1.5, and SD2.1 respectively. Similarity, Figure 20, Figure 21, and Figure 22 shows the image counts of the politicians and the imitation score of their generated images, for SD1.1, SD1.5, and SD2.1 respectively.
@@
Figure 23, Figure 24, Figure 25 show the image counts of classical artists and the similarity between their training and generated images; and Figure 26, Figure 27, Figure 28 show the image counts of modern artists and the similarity between their training and generated images. The images were generated using SD1.1, SD1.5, and SD2.1 respectively.
@@
Imitation Threshold Estimation for Human Face Imitation: In Figure 17, we observe that the imitation scores for the individuals with small image counts is close to 0 (left side), and it increases as the number of their image counts increase towards the right. The highest similarity is 0.5 and it is for the individuals in the rightmost region of the plot. The solid line in the plot shows the mean similarity over the five image generation prompt with the shaded area showing the variance over them. We observe a low variance in the imitation score among the generation prompts. And we also observe that the variance does not depend on the image counts which indicates that the performance of the face recognition model does not depend on the popularity of the individual. The change detection algorithm finds the change point to be at 364 faces for human face imitation for celebrities, when using SD1.1 for image generation. Figure 18 shows the similarity between the training and generated images when images are generated using SD1.5. Identically to SD1.1, the change is detected at 364 faces for face imitation when using SD1.5. We also performed ablation experiments with different face embeddings models and justify the choice of our model (see Appendix K). For all the plots, we also analyze the trend by using isotonic regression which learns non-decreasing linear regression weights that fits the data best.
@@
Imitation Threshold Estimation for Human Face Imitation (Politicians): Figure 20 shows the imitation scores for politicians which is very similar to the plot obtained for celebrities. We observe a low variance in the imitation score among the generation prompts. We also observe that the variance does not depend on the image counts which indicates that the performance of the face recognition model does not depend on the popularity of the individual. The change detection algorithm finds the change point to be at 234 faces for human face imitation for politicians, when using SD1.1 for image generation. Figure 21 shows the similarity between the training and generated images when images are generated using SD1.5. Similar to SD1.1, the change is detected at 234 faces.
@@
Imitation Threshold Estimation for Art Style Imitation: In Figure 23, we observe that the imitation scores for artists with low image counts have a baseline value around 0.2 (left side), and it increases as the number of their image counts increase towards the right. The highest similarity is 0.76 and it is for the artists in the rightmost region of the plot. We also observe a low variance across the generation prompts, and the variance does not depend on the image frequency of the artist. The change detection algorithm finds the change point to be at 112 images for art style imitation of classical artists, when using SD1.1 for image generation. Figure 24 shows the similarity between the training and generated images when images are generated using SD1.5. Similar to SD1.1, the change is detected at 112 faces for art style imitation when using SD1.5. These thresholds are slightly higher for style imitation of modern artists, 198 for both SD1.1 and SD1.5.
Appendix J Examples of Outliers
Figure 29 and Figure 30 show examples of outliers of the first kind, where aliases of a celebrity leads to under counting of their images in the pretraining data.
@@
Figure 31 and Figure 32 show examples of outliers of the first kind for artists, where aliases of an artist leads to under counting of their art works in the pretraining data.
@@
Appendix K Ablation Experiment with Different Face Embedding Models
@@
In this section, we show the difference in the performance of several face embedding models and justify the choice of the final choice of our face embedding model. Face embedding models are evaluated using two main metrics: false-match rate (FMR) and true-match rate (TMR) (NIST, 2020). FMR measures how many times does a model says two people are the same when they are not and TMR measures how many times a model says two people are the same when they are the same. Ideally, a face embedding model should have low FMR and high TMR. An important variant of these metrics is the disparity of FMR and TMR of a model across different demographic groups. Ideally, a model should have low disparity in these metrics across different demographics. We also focus on the variance of these metrics across demographics in making the final choice.
We evaluate the FMR and TMR of eight different face embedding models (seven open-sourced and one proprietary). The open-source models were chosen based on their popularity on Github (Serengil & Ozpinar, 2020; Deng et al., 2022; 2020), and we also experiment with Amazon Rekognition, a proprietary model. For evaluating the disparity of these metrics across different demographic groups we grouped celebrities in six demographic groups primarily categorized according to skin color tone (black, brown, and white) and perceived gender (male and female; for simplicity). Each of the six groups had 10 celebrities (a total of 60), with no intersection between them. The categorization was done manually by looking at the reference images of the celebrities. For each celebrity, we collect 10 reference images from the internet by using the procedure described in Appendix D. We use these images to compare the FMR and TMR of the face recognition models, as these images are the gold standard images of a person.
@@
Figure 33 and Figure 34 shows the FMR and TMR for six demographic groups for all the face embedding models. All the open-sourced models, except InsightFace, either have a high disparity in FMR values across the demographic groups (ArcFace, Facenet, Facenet512, DeepFace) or have very low TMR (GhostFaceNet_W1, GhostFaceNet_V1). We choose InsightFace for our experiments because of it has 1) a low overall FMR, 2) decent TMR, 3) a low disparity of FMR and TMR across the demographic groups, and 4) is open-sourced. Having a low disparity of the metrics across individuals of different demographic groups is crucial for an accurate estimation of the imitation threshold. The Amazon Rekognition model would also be a viable choice based on these metrics, however, it is not open-sourced and therefore expensive for our experiments.
Appendix L Count Distribution and the List of Sampled Entities for Each Domain
@@
We collect celebrities from https://www.popsugar.com/Celebrities and https://celebanswers.com/celebrity-list/. The distribution of the caption counts of the sampled celebrities is displayed in Table 8. The sampled celebrities in the descending order of their number of caption counts are:

### v2

@@ E.1 Collection of Reference Images for Human Faces Domain
    Candidate set: First, we retrieved the first hundred images by searching a person’s name on Google Images. We used SerpAPI (serpapi) as a wrapper perform the searches.
@@
    One strategy is for the faces to form a graph where the vertices are the face embeddings and the edges connecting two embeddings have a weight equal to the cosine similarity between them, and we select a dense k-subgraph (dense-k-subgraph-reference) from this graph. Selecting such a subgraph means finding a mutually homogeneous subset. We can find the vertices of this dense k-subgraph by cardinality-constrained submodular function minimization (bilmes2022submodularity; nagano2011size) on a facility location function (bilmes2022submodularity). We run this minimization and select a subset of images (at least of size ten) that has the highest average cosine similarity between each pair of images.
@@
[…]
@@
[…]
@@
Appendix F Implementation Details of MIMETIC2 for Human Face Imitation
F.1 Filtering of Training Images
Images whose captions mention the concept of interest often do not contain it (as shown with Mary-Lee Pfeiffer in Figure˜3). As such, we filter images where the concept does not appear in the image, which we detect using a dedicated classifier. In what follows we describe the filtering mechanism.
@@ F.1 Filtering of Training Images
We collect reference images for each person using SerpAPI as described in Appendix˜E. These images are the gold standard images that we manually vet to ensure that they contain the target person of interest (see Appendix˜E for the details). We use the reference images to filter out the images in the pretraining dataset that are not of this person. Concretely, for each person we use a face embedding model (deng2020-insightface4) to measure the similarity between the faces in the reference images and the faces in the images from the pretraining datasets whose captions mention this person. If the similarity of a face in the pretraining images to any of the faces in the reference images is above a certain threshold, that face is considered to belong to the person of interest. We determine this threshold to distinguish faces of the same person from faces of different persons in the next paragraph. Note that this procedure already filter outs any image that does not contain a face, because the face embedding model would only embed an image if it detects a face in that image.
@@
Determining Filtering Threshold: The next step is to determine the threshold for which we consider two faces to belong to the same person. For this purpose, we measure the similarity between pairs of faces of the same person and the similarity between pairs of faces of different persons Since the reference images for each person is manually vetted to be correct, we use these images for this procedure. We plot the histogram of the average similarity between the faces of the same person (blue colored) and the similarity between faces of different persons (red colored) in Figure˜10. We see that the two histograms are well separated, with the lowest similarity value between the faces of the same person being 0.56 and the highest similarity value between the faces of different persons being 0.36. Therefore, any threshold value between 0.36 and 0.56 can separate two faces of the same person, from the faces of different people. In our experiments, we use the midpoint threshold of 0.46 (true positive rate (tpr) of 100%; false positive rate (fpr) of 0%) to filter any face in the pretraining images that do not belong to the person of interest (see Table˜8 for the classifier metrics). The filtering process gives us both the image frequency a person in the pretraining data, and the pretraining images that we compare the faces in the generated images to measure the imitation score.
Table 8: Metrics for the classifier used to distinguish between the faces of individuals and the art style of artists.
Domain Accuracy (%) Precision (%) Recall (%)
Celebrity 100 100 100
Politicians 100 100 100
Classical Artists 95.0 92.3 98.3
Modern Artists 94.4 92.9 96.1
F.2 Measurement of Imitation Score
@@
To measure the imitation between the training and generated images of a person, we compute the cosine similarity between the face embeddings of the faces in their generated images and their filtered training images from the previous step. However, measuring the similarity using all the pretraining images can underestimate the actual imitation. This is because several individuals have significant variations in their faces in the pretraining images and the text-to-image model does not capture all these variations. For example, consider the pretraining images of Samuel L. Jackson in Figure˜11(a). These images have significant variations in beard, hair, and age. However, when the text-to-image model is prompted to generate images of Samuel L. Jackson, the generated images in Figure˜11(b) only show a specific facial characteristic of him (middle-aged, bald, with no or little beard). Since MIMETIC2’s goal is not to measure if a text-to-image model captures all the variations of a person, we want to reward the model even if it has only captured a particular characteristic (which it has in this case of Samuel L. Jackson). Therefore, instead of comparing the similarity of generated images to all the training images, we compare the similarity to only the ten training images that have the highest cosine similarity to the generated images on average.
Appendix G Implementation Details of MIMETIC2 for Art Style Imitation
G.1 Filtering of Training Images
@@
Filtering Non-Art Images: To filter non-art images from the pretraining dataset, we use a classifier that separates art images from non-art images. Concretely, we embed the pretraining images using a CLIP ViT-H/14 (open-clip-1) image encoder and measure the cosine similarity of the image embeddings and the text embeddings of the string ‘an artwork’, embedded using the text encoder of the same model. Only when the similarity between the embeddings is higher than a threshold described below, we consider those pretraining images as an artwork. To determine this threshold, we choose a similarity score that separates art images from non-art images. We use the images from the Wikiarts dataset (wikiarts-dataset) as the (positive) art images and MS COCO dataset images (mscoco-dataset) as the (negative) non-art images. Note that MS COCO dataset was collected by photographing everyday objects that art was not part of, making it a valid set of negative examples of art.
We plot the histogram of cosine similarity of the embeddings of art and non-art images to the text embedding of ‘an artwork’ (see Figures˜12(a) and 12(b). We observe that the art and non-art images both the artist groups are well separated (although not perfect, Figure˜13 and Figure˜14 shows examples of misclassified and correctly classfied images from both datasets). We choose the threshold that maximizes the F1 score of the separation (0.182 for the classical artists and 0.177 for the modern artists).
@@
Figure 13: Images that are misclassified by our art vs. non-art threshold in Figure˜12(a).
@@
Figure 14: Images that are correctly classified by our art vs. non-art threshold in Figure˜12(a).
@@
Filtering Images of Other Art Styles: Similar to the case of human faces, not all art images whose captions mention an artist were created by that artist. We want to filter out such images. For this purpose, we collect reference images for each artist (see Appendix˜E for details) and use them to classify the training images that belong to the artist of interest. Concretely, we measure the similarity between the pretraining images and the reference images of each artist, and only retain images whose similarity to the reference images is higher than a threshold.
To determine this threshold, we measure the similarity between pairs of art images of the same artists and pairs of art images from different artists. We embed the images using an art style embedding model (somepalli2024-style-similarity) and plot the histogram of similarities between art images of the same artist (blue colored) and art images of different artists (red colored) in Figure˜15(a) for classical artists and Figure˜15(b) for modern artists. We see that the two histograms are well separated (although not perfect, Figure˜16 shows paintings by two artists whose art style is very similar and cannot be distinguished by our threshold). We choose the threshold that maximizes the F1 score of the separation between these two groups (0.278 for classical artists and 0.288 for modern artists). See Table˜8 for the classifier metrics. The retained images give us both the image counts of each artist and the training images that we compare to the generated images to measure the imitation score.
@@ G.2 Similarity Measurement
We embed all the generated images and the filtered pretraining images using the art style embedding model (somepalli2024-style-similarity) and measure the cosine similarity between each pair of generated and pretraining images. Similar to the case of the human faces, we do not want to underestimate the art style similarity between the generated and training images by comparing the generated images to all the training images of this artist. Therefore, we measure the similarity of generated images to the ten training images that are on average the most similar to the generated images.
Appendix H Evaluation of the Imitation Thresholds
Evaluating the accuracy of the imitation thresholds MIMETIC2 finds is a very expensive task, it requires conducting the optimal experiment described in Section˜3 (training $\mathcal{O}(\log m)$ models with different number of images of each concept). In order to make the evaluation tractable, we finetune a pretrained text-to-image model with concepts that the model has not seen during pretraining and find the number of images required for the model to be able to imitate them. We find that all the concepts require more than 250 images for imitation when the models are finetuned on them, which is very close to the imitation threshold MIMETIC2 finds for this pretrained model without finetuning its (234 images). Based on these results, we conclude that the imitation thresholds estimated by MIMETIC2 are accurate.
We do this by searching politicians who became popular after the LAION-2B-en dataset was released (after 2022) (§C.2) and collecting their images from Google Images. We finetune SD1.4 on images of these politicians, starting with 50 images of a politician (much below the imitation threshold found by MIMETIC2) and going up to 800 images (above the estimated imitation threshold). We finetune the model separately for each politician. To mimic the original training setup of the SD models (where images of a concept are naturally interspersed with other images in the dataset), we mix the politician images with other images taken randomly from the LAION-2B-en dataset such that for each finetuning experiment the total number of images added upto 10,000 (9,950 random images mixed with 50 images of the politician, 9,900 random images mixed with 100 images of the politician, and so on for each finetuning experiment). We finetune on this new dataset for a single epoch with a learning rate of 5e55e-5.
We want to find the number of images required for imitation to occur in this setup. Imitation can be established either with a human study or an automatic metric that correlates with human perception of imitation. Since we already conducted a human study and found that our imitation threshold correlates with human perception, we argue that the automatic imitation score for concepts that are on the right hand side of the imitation threshold would also correlate with the human perception of imitation. MIMETIC2 estimated that the imitation score for concepts on the right hand side of the imitation threshold for face imitation of politicians for SD1.4 is 0.26 (Figure˜19). We deem the number of images required to reach this imitation score as the number of images required for imitation for this finetuning experiment.
We report the results in Table˜9. It shows that the imitation score remains lower than 0.26 when less than 200 images are used to finetune the SD model and the score reaches 0.26 only when finetuned with 250 or more images of a politician.
Thus the number of images required to reach the imitation score of 0.26 in the finetuning experiment (250 images) is very close to the imitation threshold MIMETIC2 finds for SD1.4 without requiring to finetune the model (234 images). Based on these results, we conclude that the imitation thresholds estimated by MIMETIC2 are accurate.
Table 9: The imitation scores of the concepts as SD1.4 is finetuned on increasing number of images of a concept. We report imitation scores averaged over 10 such politicians. We note that the imitation scores have values 0.26 or greater when the model is finetuned on 250 or more images of a politician. Critically, this is the average imitation score for politicians on the right hand side of the imitation threshold in Figure˜19 indicating that 250 images or more are required for imitation in this experiment.
# Finetuning Images 50 100 150 200 250 300 400 500 600 700 800
Avg. Imitation Score 0.07 0.16 0.19 0.21 0.26 0.27 0.29 0.31 0.32 0.36 0.34
Appendix I Imitation Thresholds of SD models in Series 1 and 2
@@
Our experimental results in Section˜6 found that for most domains the imitation thresholds for SD1.1 and SD1.5 are almost the same, while being higher for SD2.1. We hypothesized that the difference is due to their different text encoders. All models in SD1 series use the same text encoder from CLIP, whereas SD2.1 uses the text encoder from OpenCLIP. To test the validity of this hypothesis, we repeated the experiments for all models in SD1 series for politicians and computed their imitation thresholds. Table˜10 shows the thresholds for the politicians. We find that the imitation thresholds for all the models in SD1 series is almost the same, and is lower than the threshold for SD2.1 model. This evidence supports our hypothesis of the difference in the text-encoders being the main reason for the difference in the imitation thresholds.
@@
Appendix J Change Points
Table˜11 we show all the change points that PELT found for each experiment (Table˜3 reports the first change point as the imitation threshold).
Table 11: Imitation Thresholds for human face and art style imitation for the different text-to-image models and datasets we experiment with. This table shows all the change points that PELT found for each experiment (Table˜3 reports the first change point as the imitation threshold).
@@
Appendix K All Results: The Imitation Threshold
Figure 20: Human Face Imitation (Celebrities): Similarity between the training and generated images for all celebrities. The celebrities with zero image counts are shaded with light gray. We show the mean and variance over the five generation prompts. The images were generated using LDM. The change point for human face imitation for celebrities when generating images using LDM is detected at 648 faces.
@@
Figure 24: Human Face Imitation (Politicians): Similarity between the training and generated images for all politicians. The politicians with zero image counts are shaded with light gray. We show the mean and variance over the five generation prompts. The images were generated using LDM. The change point for human face imitation for politicians when generating images using LDM is detected at 309 faces.
@@
In this section, we estimate the imitation threshold for human face and art style imitation for three different text-to-image models. Figure˜20, Figure˜21, Figure˜22, and Figure˜23 show the image counts of celebrities on the x-axis (sorted in increasing order of image counts) and the imitation score of their generated images (averaged over the five image generation prompts) on the y-axis. The images were generated using LDM, SD1.1, SD1.5, and SD2.1 respectively. Similarity, Figure˜24, Figure˜25, Figure˜26, and Figure˜27 shows the image counts of the politicians and the imitation score of their generated images, for LDM, SD1.1, SD1.5, and SD2.1 respectively.
@@
Figure˜28, Figure˜29, Figure˜30, Figure˜31 show the image counts of classical artists and the similarity between their training and generated images; and Figure˜32, Figure˜33, Figure˜34, Figure˜35 show the image counts of modern artists and the similarity between their training and generated images. The images were generated using LDM, SD1.1, SD1.5, and SD2.1 respectively.
@@
Imitation Threshold Estimation for Human Face Imitation: In Figure˜21, we observe that the imitation scores for the individuals with small image counts is close to 0 (left side), and it increases as the number of their image counts increase towards the right. The highest similarity is 0.5 and it is for the individuals in the rightmost region of the plot. The solid line in the plot shows the mean similarity over the five image generation prompt with the shaded area showing the variance over them. We observe a low variance in the imitation score among the generation prompts. And we also observe that the variance does not depend on the image counts which indicates that the performance of the face recognition model does not depend on the popularity of the individual. The change detection algorithm finds the change point to be at 364 faces for human face imitation for celebrities, when using SD1.1 for image generation. Figure˜22 shows the similarity between the training and generated images when images are generated using SD1.5. Identically to SD1.1, the change is detected at 364 faces for face imitation when using SD1.5. We also performed ablation experiments with different face embeddings models and justify the choice of our model (see Appendix˜M). For all the plots, we also analyze the trend by using isotonic regression which learns non-decreasing linear regression weights that fits the data best.
@@
Imitation Threshold Estimation for Human Face Imitation (Politicians): Figure˜25 shows the imitation scores for politicians which is very similar to the plot obtained for celebrities. We observe a low variance in the imitation score among the generation prompts. We also observe that the variance does not depend on the image counts which indicates that the performance of the face recognition model does not depend on the popularity of the individual. The change detection algorithm finds the change point to be at 234 faces for human face imitation for politicians, when using SD1.1 for image generation. Figure˜26 shows the similarity between the training and generated images when images are generated using SD1.5. Similar to SD1.1, the change is detected at 234 faces.
Figure 28: Art Style Imitation (Classical Artists): similarity between the training and generated images for classical art styles. We show the mean and variance over the five generation prompts. The images were generated using LDM. The change point for art style imitation when generating images using LDM is detected at 312 images.
@@
Figure 32: Art Style Imitation (Modern Artists): similarity between the training and generated images for modern art styles. We show the mean and variance over the five generation prompts. The images were generated using LDM. The change point for art style imitation when generating images using SD0 is detected at 282 images.
@@
Imitation Threshold Estimation for Art Style Imitation: In Figure˜29, we observe that the imitation scores for artists with low image counts have a baseline value around 0.2 (left side), and it increases as the number of their image counts increase towards the right. The highest similarity is 0.76 and it is for the artists in the rightmost region of the plot. We also observe a low variance across the generation prompts, and the variance does not depend on the image frequency of the artist. The change detection algorithm finds the change point to be at 112 images for art style imitation of classical artists, when using SD1.1 for image generation. Figure˜30 shows the similarity between the training and generated images when images are generated using SD1.5. Similar to SD1.1, the change is detected at 112 faces for art style imitation when using SD1.5. These thresholds are slightly higher for style imitation of modern artists, 198 for both SD1.1 and SD1.5.
Appendix L Examples of Outliers
Figure˜36 and Figure˜37 show examples of outliers of the first kind, where aliases of a celebrity leads to under counting of their images in the pretraining data.
@@
Figure˜38 and Figure˜39 show examples of outliers of the first kind for artists, where aliases of an artist leads to under counting of their art works in the pretraining data.
@@
Appendix M Ablation Experiment with Different Face Embedding Models
@@
In this section, we show the difference in the performance of several face embedding models and justify the choice of the final choice of our face embedding model. Face embedding models are evaluated using two main metrics: false-match rate (FMR) and true-match rate (TMR) (nist-face-recognition-test). FMR measures how many times does a model says two people are the same when they are not and TMR measures how many times a model says two people are the same when they are the same. Ideally, a face embedding model should have low FMR and high TMR. An important variant of these metrics is the disparity of FMR and TMR of a model across different demographic groups. Ideally, a model should have low disparity in these metrics across different demographics. We also focus on the variance of these metrics across demographics in making the final choice.
We evaluate the FMR and TMR of eight different face embedding models (seven open-sourced and one proprietary). The open-source models were chosen based on their popularity on Github (serengil2020-deepface; deng2018-insightface3; deng2020-insightface4), and we also experiment with Amazon Rekognition, a proprietary model. For evaluating the disparity of these metrics across different demographic groups we grouped celebrities in six demographic groups primarily categorized according to skin color tone (black, brown, and white) and perceived gender (male and female; for simplicity). Each of the six groups had 10 celebrities (a total of 60), with no intersection between them. The categorization was done manually by looking at the reference images of the celebrities. For each celebrity, we collect 10 reference images from the internet by using the procedure described in Appendix˜E. We use these images to compare the FMR and TMR of the face recognition models, as these images are the gold standard images of a person.
@@
Figure˜40 and Figure˜41 shows the FMR and TMR for six demographic groups for all the face embedding models. All the open-sourced models, except InsightFace, either have a high disparity in FMR values across the demographic groups (ArcFace, Facenet, Facenet512, DeepFace) or have very low TMR (GhostFaceNet_W1, GhostFaceNet_V1). We choose InsightFace for our experiments because of it has 1) a low overall FMR, 2) decent TMR, 3) a low disparity of FMR and TMR across the demographic groups, and 4) is open-sourced. Having a low disparity of the metrics across individuals of different demographic groups is crucial for an accurate estimation of the imitation threshold. The Amazon Rekognition model would also be a viable choice based on these metrics, however, it is not open-sourced and therefore expensive for our experiments.
Appendix N Count Distribution and the List of Sampled Entities for Each Domain
@@
We collect celebrities from https://www.popsugar.com/Celebrities and https://celebanswers.com/celebrity-list/. The distribution of the caption counts of the sampled celebrities is displayed in Table˜12. The sampled celebrities in the descending order of their number of caption counts are: